OpenStudy (anonymous):

If sinx+cosx=a then what is sin^3 x +cos^3 x in terms of "a"?

OpenStudy (loser66):

ha,!! I can imitate oldrinbataku to solve this problem.

OpenStudy (loser66):

sin^3 + cos^3 = (sin + cos) (sin^2 -sincos +cos^2) = (sin + cos) ( 1-sin cos)

OpenStudy (loser66):

and (sin + cos)^2 = sin^2 + cos ^2 + 2sin cos = 1 + 2 sincos = a^2 so, sin cos = (a^2 -1)/2 replace all (sin + cos) (1-sin cos) = a (1 - (a^2 -1) /2) let write it in neat \[sin^3x+cos^3 x= a(1-\dfrac{a^2-1}{2})\\\] =\(\dfrac{2a-a^2+1}{2}\)

OpenStudy (loser66):

@oldrin.bataku check your mentee's work, hihihi

OpenStudy (anonymous):